# Modeling Spatio-Temporal Transport: From Rigid Advection to Realistic Dynamics

### v3

@@ Abstract
Stochastic models for spatio-temporal transport face a critical trade-off between physical realism and interpretability. The advection model with a single constant velocity is interpretable but physically limited by its perfect correlation over time. This work aims to bridge the gap between this simple framework and its physically realistic extensions. Our guiding principle is to introduce a spatial correlation structure that vanishes over time. To achieve this, we present two distinct approaches. The first constructs complex velocity structures, either through superpositions of advection components or by allowing the velocity to vary locally. The second is a spectral technique that replaces the singular spectrum of rigid advection with a more flexible form, introducing temporal decorrelation controlled by parameters. We accompany these models with efficient simulation algorithms and demonstrate their success in replicating complex dynamics, such as tropical cyclones and the solutions of partial differential equations.
@@ 1 Introduction
To model these dynamics, one major research direction uses Lagrangian approaches to build non-stationary covariance structures by tracking particle trajectories. This framework has proven effective for modeling environmental processes like ocean waves and precipitation (ailliot2011; benoit2018), as these phenomena are inherently governed by physical transport. Another line of research employs Stochastic Partial Differential Equations (SPDEs) to embed physical principles like advection and diffusion directly into the models. This has led to computationally efficient models with interpretable parameters (sigrist2015), which have since been extended to handle more complex, spatially-varying dynamics (liu2022; carlotto2024). While other methods focus on simulating broader classes of non-separable covariance functions (see for instance allard2020simulating), they often face a trade-off: they either omit explicit directional transport for simplicity, or incorporate it at the cost of significant model complexity. Our work seeks to occupy a middle ground, prioritizing both physical realism and interpretability. For comprehensive reviews on the state of dynamic and non-separable models, a topic that builds on foundational work like gneiting2002nonseparable, porcu2020nonseparable, chen2021space, and papalexiou2021advancing.
@@
Thus, this paper makes a threefold contribution. First, we introduce these novel classes of models for advection. Second, we provide computationally efficient algorithms to simulate them, ensuring that these methods are accessible for practical application. We demonstrate that these algorithms are competitive with, and in some cases superior to, standard approaches by avoiding known issues like numerical diffusion. Third, we demonstrate the applicability of these models by demonstrating their ability to generate complex, physically plausible dynamics, such as those of a tropical cyclone, and to replicate the solutions of well-known transport and diffusion PDEs. To facilitate their use by practitioners, we provide implementations of these methods in R (Rmanual).
The paper is organized as follows. Section 2 formally defines the frozen field model and discusses two primary simulation approaches. Section 3 introduces our first class of generalizations based on velocity structures, presents simulated examples, and compares the results to numerical PDE solvers. Section 4 details our second generalization strategy based on spectral modifications and analyzes the role of the model parameters. Finally, Section 5 provides concluding remarks and outlines directions for future research. Further results are included in the supplementary materials.
@@ 2.1 Framework and definition
A common and important simplifying assumption is spatial stationarity. A spatial random field $X_{S}(\mathbf{s})$ is said to be stationary if its mean is constant (assumed to be zero without loss of generality for covariance modeling) and its covariance function $c_{XX}(\mathbf{s}_{1},\mathbf{s}_{2})$ depends only on the lag $\mathbf{h}=|\mathbf{s}_{2}-\mathbf{s}_{1}|\in\mathbb{R}^{2}$. With a slight abuse of notation, we write this stationary covariance function as
@@
where $\nabla Z(\mathbf{s},t)=\left(\frac{\partial Z(\mathbf{s},t)}{\partial s_{1}}\frac{\partial Z(\mathbf{s},t)}{\partial s_{2}}\right)^{\top}$. Moreover, such process enjoys the following second order properties:
@@
$S_{ZZ}(\mathbf{k},\omega)$ $=S_{XX}(\mathbf{k})\cdot\delta(\omega+\mathbf{k}^{T}\mathbf{v}),$ (4)
where $(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}$ denotes the pair of wavenumber and frequency, respectively. Specifically, covariance $c_{ZZ}(\mathbf{h},\tau)$ consists of the covariance function $c_{XX}(\cdot)$ of the space random field evaluated at spatial lag $|\mathbf{h}-\mathbf{v}\tau|$, and it is clearly not separable. The spectrum $S_{ZZ}(\mathbf{k},\omega)$ depends on the spectrum of the space random field $S_{XX}(\mathbf{k})$, which is multiplied by a generalized function corresponding to a Dirac delta depending on velocity $\mathbf{v}$.
@@ 2.2 Simulation of Frozen Fields
Firstly, a FF can be generated by simulating a single, large spatial random field, and then advecting $\mathcal{G}$ across this field over time. More precisely, we first need to generate the extended spatial field $X_{S}^{\text{ext}}(\mathbf{s})$ on grid $\mathcal{G}^{*}=[0,N]\times[0,N]$, with $N>n$. This initial simulation can be performed with circulant embedding (ChanWood97; DietrichNewsam97; Graham2018), a non-exact spectral method relying on the use of Discrete Fourier Transforms (DFTs) and their implementation on regular grids, called Fast Fourier Transforms (FFTs). The computational cost of circulant embedding, for which a FFT is carried out once, is $\mathcal{O}(N^{2}\log N)$. Notice that the simulation domain $\mathcal{G}^{*}$ must be large enough to contain the full trajectory of the smaller observation grid $\mathcal{G}$. The required grid size $N$ should be chosen based on the properties of the spatio-temporal field being simulated. Accordingly, we set $N=n+Tv_{\text{max}}$, where $v_{\text{max}}=\max\{|v_{1}|,|v_{2}|\}$ is the maximum component of the constant velocity $\mathbf{v}=(v_{1},v_{2})^{T}$.
@@
A comparison of computational costs reveals a clear trade-off between this shifting method and a direct space-time simulation via 3D circulant embedding, which costs $\mathcal{O}(n^{2}T\log(nT))$. The total cost of the shifting method, $\mathcal{O}(N^{2}\log N+n^{2}T)$, depends heavily on $N$, which, by construction, grows linearly with $T$. Consequently, the shifting method is significantly more efficient when the number of time steps $T$ is small relative to the spatial grid dimension $n$. For long large $T$, the 3D spectral approach becomes more advantageous as its cost scales more gracefully with time. For context, both methods are vastly more efficient than an exact simulation using Cholesky decomposition of the full space-time covariance matrix, which would require a cost of $\mathcal{O}(n^{6}T^{3})$. The overall procedure for the shifting method is detailed in Algorithm 1.
@@
A second approach is an adaptation of the spectral method called Davies-Harte algorithm (see for instance craigmile2003). This method simulates the Gaussian random field based on its spectral density function. The procedure, detailed in Algorithm 2, begins by defining padded grid dimensions, $n_{\text{pad}}=2n$ and $T_{\text{pad}}=2T$. This step is crucial for mitigating the periodic boundary effects that are inherent in any FFT-based simulation.
@@
3:Define wavenumber vector $\mathbf{k}_{\text{vec}}$ for a grid of size $n_{\text{pad}}$
4:Define frequency vector $\omega_{\text{vec}}$ for a grid of size $T_{\text{pad}}$
@@ 3.1 Framework
Notice that considering a random velocity in equation (1) is an existing extension to the FF model. For instance, schlather2010some generates a Gaussian velocity with mean $\boldsymbol{\mu}\in\mathbb{R}^{2}$ and covariance $D\in\mathbb{R}^{2\times 2}$, i.e., $\mathbf{v}\sim\mathcal{N}(\boldsymbol{\mu},D)$, once for the entire field, while christakos2017spatiotemporal considers $\mathbf{v}^{t}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\boldsymbol{\mu},D)$, where the velocity $\mathbf{v}^{t}$ is drawn independently at every time point $t$. With the Distributed FF model, we offer a more complex alternative by allowing for random geometric transformations of the velocity, enabling multiple propagation directions and scales within the same field. The diminished correlation in the distributed case arises from the superposition of multiple transformed components introducing variability in propagation directions and scales.
@@
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))$ $=\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{j}\mathbf{v})\,c_{XX}\left(\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})\tau-(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t\right),$ (7)
@@
$\left((\mathbf{M}_{j}\mathbf{v})\tau+(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t,\tau\right),\quad i,j=1,\dots,n_{\text{v}}.$ (8)
@@
$=\mathrm{cov}\left(\sum_{i=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})X_{S}\left(\mathbf{s}-(\mathbf{M}_{i}\mathbf{v})t\right),\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{j}\mathbf{v})X_{S}\left(\mathbf{s}+\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})p(\mathbf{M}_{j}\mathbf{v})\,\mathrm{cov}\left(X_{S}\left(\mathbf{s}-(\mathbf{M}_{i}\mathbf{v})t\right),X_{S}\left(\mathbf{s}+\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p(\mathbf{M}_{i}\mathbf{v})p(\mathbf{M}_{j}\mathbf{v})\,c_{XX}\left(\mathbf{h}-(\mathbf{M}_{j}\mathbf{v})\tau-(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t\right).$
@@
$\mathbf{h}=(\mathbf{M}_{j}\mathbf{v})\tau+(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t,\quad i,j=1,\dots,n_{\text{v}},$
@@
This proposition reveals how the Distributed FF model generalizes the classic FF by examining the structure of its covariance in (7). First, consider the terms where $i=j$. In this case, the expression simplifies to that of a standard FF: the time-dependent component $(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t$ vanishes from the covariance argument, and the propagation path (8) reduces to simple, straight-line advection, $\left((\mathbf{M}_{i}\mathbf{v})\tau,\tau\right)$. These terms represent a weighted mixture of independent FFs, each with its own transformed velocity. Conversely, the terms where $i\neq j$ introduce novel cross-covariances between the different components. Specifically, $(\mathbf{M}_{j}\mathbf{v}-\mathbf{M}_{i}\mathbf{v})t$ denotes the growing spatial displacement between two component paths over time $t$. The presence of this term makes the overall covariance function dependent on time, thus making the Distributed FF model non-stationary. This complex behavior is known in physics as interference, referring to the interactions introduced by the superimposition of waves traveling in the same medium. Consequently, the distributed frozen field satisfies $\rho(v\tau,\tau)<1$ unless $\mathbf{M}_{i}\mathbf{v}=\mathbf{v}$ for all $i$, thereby preventing unrealistic perfect correlation along the propagation path of the FF.
@@
A random field $Z(\mathbf{s},t)$ is a Evolving Frozen Field if there is a deterministic velocity field $\mathbf{v}(\mathbf{s},t)\in{\mathbb{R}}^{2}$, and a random field $X_{S}(\mathbf{s})$ with $\mathbf{s}\in{\mathbb{R}}^{2}$ and covariance function $c_{XX}(\mathbf{h})$ such that
@@
Hence, model (9) is highly relevant, and determining its mathematical properties is intrinsecally linked to the unobserved velocity field $\mathbf{v}(\mathbf{s},t)$.
@@
with ${\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)=\begin{pmatrix}\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{2}}\\ \frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{2}}\end{pmatrix}$ and $D_{t}(\mathbf{s},t)=\left(\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial t},\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial t}\right)^{T}$. Moreover, the propagation path of $Z(\mathbf{s},t)$ is
@@
$\mathrm{cov}\left\{X(\mathbf{s}-\mathbf{v}(\mathbf{s},t)t),X(\mathbf{s}+\mathbf{h}-\mathbf{v}(\mathbf{s}+\mathbf{h},t+\tau)(t+\tau))\right\}=$

### v4

@@ Abstract
Stochastic models for spatio-temporal transport face a critical trade-off between physical realism and interpretability. The advection model with a single constant velocity is interpretable but physically limited by its perfect correlation over time. This work aims to bridge the gap between this simple framework and its physically realistic extensions. Our guiding principle is to introduce a spatial correlation structure that vanishes over time. To achieve this, we present two distinct approaches. The first constructs complex velocity structures, either through superpositions of advection components or by allowing the velocity to vary locally. The second is a spectral technique that replaces the singular spectrum of rigid advection with a more flexible form, introducing temporal decorrelation controlled by parameters. We accompany these models with efficient simulation algorithms and demonstrate their success in replicating complex dynamics, such as tropical cyclones and the solutions of partial differential equations. Finally, we illustrate the practical utility of the proposed framework by comparing its simulations to real-world precipitation data from Hurricane Florence.
@@ 1 Introduction
To model these dynamics, one major research direction uses Lagrangian approaches to build non-stationary covariance structures by tracking particle trajectories. This framework has proven effective for modeling environmental processes like ocean waves and precipitation (ailliot2011; benoit2018), as these phenomena are inherently governed by physical transport. Another line of research employs Stochastic Partial Differential Equations (SPDEs) to embed physical principles like advection and diffusion directly into the models. This has led to computationally efficient models with interpretable parameters (sigrist2015), which have since been extended to handle more complex, spatially-varying dynamics (liu2022; carlotto2024). While other methods focus on simulating broader classes of non-separable covariance functions (see for instance allard2020simulating), they often face a trade-off: they either omit explicit directional transport for simplicity, or incorporate it at the cost of significant model complexity. Our work seeks to occupy a middle ground, prioritizing both physical realism and interpretability. For comprehensive reviews on the state of dynamic and non-separable models, namely models that allow spatial and temporal dependence to interact, we refer to gneiting2002nonseparable, porcu2020nonseparable, chen2021space, and papalexiou2021advancing.
@@
Thus, this paper makes a threefold contribution. First, we introduce these novel classes of models for advection. Second, we provide computationally efficient algorithms to simulate them, ensuring that these methods are accessible for practical application. We demonstrate that these algorithms are competitive with, and in some cases superior to, standard approaches by avoiding known issues like numerical diffusion, namely the artificial spreading caused by discretization schemes. Third, we showcase the applicability of these models by illustrating their ability to generate complex, physically plausible dynamics, such as those of a tropical cyclone, and to replicate the solutions of well-known transport and diffusion PDEs. We further validate our approach with a real-data application to Hurricane Florence, comparing our simulations directly to observational precipitation records. To facilitate their use by practitioners, we provide implementations of these methods in R (Rmanual).
The paper is organized as follows. Section 2 formally defines the frozen field model and discusses two primary simulation approaches. Section 3 introduces our first class of generalizations based on velocity structures, presents simulated examples, compares the results to numerical PDE solvers, and assesses the predictive performance against observed data from Hurricane Florence. Section 4 details our second generalization strategy based on spectral modifications and analyzes the role of the model parameters. Finally, Section 5 provides concluding remarks and outlines directions for future research. Further results are included in the supplementary materials.
@@ 2.1 Framework and definition
A common and important simplifying assumption is spatial stationarity. A spatial random field $X_{S}(\mathbf{s})$ is said to be stationary if its mean is constant (assumed to be zero without loss of generality for covariance modeling) and its covariance function $c_{XX}(\mathbf{s}_{1},\mathbf{s}_{2})$ depends only on the lag $\mathbf{h}=(\mathbf{s}_{2}-\mathbf{s}_{1})\in\mathbb{R}^{2}$. With a slight abuse of notation, we write this stationary covariance function as
@@
where $\nabla Z(\mathbf{s},t)=\left(\frac{\partial Z(\mathbf{s},t)}{\partial s_{1}}\frac{\partial Z(\mathbf{s},t)}{\partial s_{2}}\right)^{\top}$. Moreover, this process has the following second order properties:
@@
$S_{ZZ}(\mathbf{k},\omega)$ $=S_{XX}(\mathbf{k})\cdot\delta(\omega+\mathbf{k}^{\top}\mathbf{v}),$ (4)
where $(\mathbf{k},\omega)\in\mathbb{R}^{2}\times\mathbb{R}$ denotes the pair of wavenumber and frequency, respectively. Specifically, the covariance $c_{ZZ}(\mathbf{h},\tau)$ consists of the covariance function $c_{XX}(\cdot)$ of the space random field evaluated at spatial lag $(\mathbf{h}-\mathbf{v}\tau)$. This structure makes the FF model non-separable, as the spatial and temporal dependencies are intrinsically coupled through the advection term $\mathbf{v}\tau$. The spectrum $S_{ZZ}(\mathbf{k},\omega)$, i.e. the Fourier transform of the covariance function $c_{ZZ}(\mathbf{h},\tau)$, depends on the spectrum of the space random field $S_{XX}(\mathbf{k})$, which is multiplied by a generalized function corresponding to a Dirac delta depending on velocity $\mathbf{v}$.
@@ 2.2 Simulation of Frozen Fields
Firstly, a FF can be generated by simulating a single, large spatial random field, and then advecting $\mathcal{G}$ across this field over time. More precisely, we first need to generate the extended spatial field $X_{S}^{\text{ext}}(\mathbf{s})$ on grid $\mathcal{G}^{*}=[0,N]\times[0,N]$, with $N>n$. This initial simulation can be performed with circulant embedding (ChanWood97; DietrichNewsam97; Graham2018), a non-exact spectral method relying on the use of Discrete Fourier Transforms (DFTs) and their implementation on regular grids, called Fast Fourier Transforms (FFTs). The computational cost of circulant embedding, for which a FFT is carried out once, is $\mathcal{O}(N^{2}\log N)$. Notice that the simulation domain $\mathcal{G}^{*}$ must be large enough to contain the full trajectory of the smaller observation grid $\mathcal{G}$. The required grid size $N$ should be chosen based on the properties of the spatio-temporal field being simulated. Accordingly, we set $N=n+Tv_{\text{max}}$, where $v_{\text{max}}=\max\{|v_{1}|,|v_{2}|\}$ is the maximum component of the constant velocity $\mathbf{v}=(v_{1},v_{2})^{\top}$.
@@
A comparison of computational costs reveals a clear trade-off between this shifting method and a direct space-time simulation via 3D circulant embedding, which costs $\mathcal{O}(n^{2}T\log(nT))$. The total cost of the shifting method, $\mathcal{O}(N^{2}\log N+n^{2}T)$, depends heavily on $N$, which, by construction, grows linearly with $T$. Consequently, the shifting method is significantly more efficient when the number of time steps $T$ is small relative to the spatial grid dimension $n$. For long $T$, the 3D spectral approach becomes more advantageous as its cost scales less aggressively with time. For context, both methods are vastly more efficient than an exact simulation using Cholesky decomposition of the full space-time covariance matrix, which would require a cost of $\mathcal{O}(n^{6}T^{3})$. The overall procedure for the shifting method is detailed in Algorithm 1.
@@
A second approach is an adaptation of the spectral method called Davies-Harte algorithm (see for instance craigmile2003). This method simulates the Gaussian random field based on its spectral density function. The procedure, detailed in Algorithm 2, begins by defining padded grid dimensions, $n_{\text{pad}}=2n$ and $T_{\text{pad}}=2T$. This step is crucial for mitigating the periodic boundary effects that are inherent in any FFT-based simulation. The wavenumber and frequency vectors, $k_{vec}$ and $\omega_{vec}$, are defined as regular sequences spanning the fundamental frequency domain, with uniform spacing $1/n_{pad}$ and $1/T_{pad}$ respectively, ensuring compatibility with the discrete Fourier transform.
@@
3:Define wavenumber vector $\mathbf{k}_{\text{vec}}$ for a grid of size $n_{\text{pad}}$ $\triangleright$ Fourier frequencies
4:Define frequency vector $\omega_{\text{vec}}$ for a grid of size $T_{\text{pad}}$ $\triangleright$ Fourier frequencies
@@ 3.1 Framework
Notice that considering a random velocity in equation (1) is an existing extension to the FF model. For instance, schlather2010some considers the Cox-Isham rainfall model (cox1988) where $\mathbf{v}\sim\mathcal{N}(\boldsymbol{\mu},D)$ is drawn once for the entire field. In that framework, the mean velocity $\boldsymbol{\mu}\in\mathbb{R}^{2}$ determines the dominant advection direction, while the covariance matrix $D\in\mathbb{R}^{2\times 2}$ governs the spread of velocities around this mean. The Distributed FF provides a discrete mixture generalization of this construction. To simplify the notation, let $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$ and $p_{i}=p(\mathbf{v}_{i})$ for $i=1,\dots,n_{v}$. Then, the empirical mean and covariance of the velocities $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$ with weights $\{p_{i}\}_{i=1}^{n_{v}}$ are
$\bar{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\mathbf{v}_{i},\qquad\mathbf{C}_{v}=\sum_{i=1}^{n_{v}}p_{i}(\mathbf{v}_{i}-\bar{\mathbf{v}})(\mathbf{v}_{i}-\bar{\mathbf{v}})^{\top},$
which play the roles of $\boldsymbol{\mu}$ and $\mathbf{D}$ in the Gaussian formulation of Schlather (2010), respectively. This discrete mixture representation offers greater modeling flexibility, as it allows for arbitrary velocity distributions through the choice of transformations $\{\mathbf{M}_{i}\}_{i=1}^{n_{v}}$ and weights $\{p_{i}\}_{i=1}^{n_{v}}$.
By allowing for random geometric transformations of the velocity, the Distributed FF model enables multiple propagation directions and scales within the same field. This structure naturally diminishes temporal correlation, as the superposition of components propagating with different velocities disrupts the perfect dependence observed in the frozen field model.
@@
$\mathrm{cov}(Z(\mathbf{s},t),Z(\mathbf{s}+\mathbf{h},t+\tau))$ $=\sum_{i=1}^{n_{\text{v}}}p_{i}\sum_{j=1}^{n_{\text{v}}}p_{j}\,c_{XX}\left(\mathbf{h}-\mathbf{v}_{j}\tau-(\mathbf{v}_{j}-\mathbf{v}_{i})t\right),$ (7)
@@
$\left(\mathbf{v}_{j}\tau+(\mathbf{v}_{j}-\mathbf{v}_{i})t,\tau\right),\quad i,j=1,\dots,n_{\text{v}},$ (8)
where $\mathbf{v}_{i}=\mathbf{M}_{i}\mathbf{v}$ and $p_{i}=p(\mathbf{M}_{i}\mathbf{v})$ for $i=1,\dots,n_{v}$.
@@
$=\mathrm{cov}\left(\sum_{i=1}^{n_{\text{v}}}p_{i}X_{S}\left(\mathbf{s}-\mathbf{v}_{i}t\right),\sum_{j=1}^{n_{\text{v}}}p_{j}X_{S}\left(\mathbf{s}+\mathbf{h}-\mathbf{v}_{j}(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p_{i}p_{j}\,\mathrm{cov}\left(X_{S}\left(\mathbf{s}-\mathbf{v}_{i}t\right),X_{S}\left(\mathbf{s}+\mathbf{h}-\mathbf{v}_{j}(t+\tau)\right)\right)$
$=\sum_{i=1}^{n_{\text{v}}}\sum_{j=1}^{n_{\text{v}}}p_{i}p_{j}\,c_{XX}\left(\mathbf{h}-\mathbf{v}_{j}\tau-(\mathbf{v}_{j}-\mathbf{v}_{i})t\right).$
@@
$\mathbf{h}=\mathbf{v}_{j}\tau+(\mathbf{v}_{j}-\mathbf{v}_{i})t,\quad i,j=1,\dots,n_{\text{v}},$
@@
This proposition reveals how the Distributed FF model generalizes the classic FF by examining the structure of its covariance in (7). First, consider the terms where $i=j$. In this case, the expression simplifies to that of a standard FF: the time-dependent component $(\mathbf{v}_{j}-\mathbf{v}_{i})t$ vanishes from the covariance argument, and the propagation path (8) reduces to simple, straight-line advection, $\left(\mathbf{v}_{i}\tau,\tau\right)$. These terms represent a weighted mixture of independent FFs, each with its own transformed velocity. Conversely, the terms where $i\neq j$ introduce novel cross-covariances between the different components. Specifically, $(\mathbf{v}_{j}-\mathbf{v}_{i})t$ denotes the growing spatial displacement between two component paths over time $t$. The presence of this term makes the overall covariance function dependent on time, thus making the Distributed FF model non-stationary. This complex behavior is known in physics as interference, referring to the interactions introduced by the superimposition of waves traveling in the same medium. Consequently, the distributed frozen field satisfies $\rho(\mathbf{v}\tau,\tau)<1$ unless $\mathbf{v}_{i}=\mathbf{v}$ for all $i$, thereby preventing unrealistic perfect correlation along the propagation path of the FF.
@@
A random field $Z(\mathbf{s},t)$ is an Evolving Frozen Field if there is a deterministic velocity field $\mathbf{v}(\mathbf{s},t)\in{\mathbb{R}}^{2}$, and a random field $X_{S}(\mathbf{s})$ with $\mathbf{s}\in{\mathbb{R}}^{2}$ and covariance function $c_{XX}(\mathbf{h})$ such that
@@
Hence, model (9) is highly relevant, and determining its mathematical properties is intrinsically linked to the unobserved velocity field $\mathbf{v}(\mathbf{s},t)$.
@@
with ${\mathbf{D}}_{\mathbf{s}}(\mathbf{s},t)=\begin{pmatrix}\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial s_{2}}\\ \frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{1}}&\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial s_{2}}\end{pmatrix}$ and $D_{t}(\mathbf{s},t)=\left(\frac{\partial\text{v}_{1}(\mathbf{s},t)}{\partial t},\frac{\partial\text{v}_{2}(\mathbf{s},t)}{\partial t}\right)^{\top}$. Moreover, the propagation path of $Z(\mathbf{s},t)$ is
@@
$\mathrm{cov}\left(X(\mathbf{s}-\mathbf{v}(\mathbf{s},t)t),X(\mathbf{s}+\mathbf{h}-\mathbf{v}(\mathbf{s}+\mathbf{h},t+\tau)(t+\tau))\right)=$